OpenStudy (llamasaremine):

I Need some help i have been doing this forever typing to figure it out. will Fan and Medal

OpenStudy (faiqraees):



OpenStudy (llamasaremine):

thats what i have to do. sorry thought i clicked the whole picture.

OpenStudy (faiqraees):

Okay start by calculating the length of the base of the triangle with angle 36

OpenStudy (llamasaremine):

idk how to do that..

OpenStudy (faiqraees):

Hint: Trigonometric Ratios

OpenStudy (anonymous):

sorry but i think it in this way 2x-12=36 2x = 48 x = 48/2 x = 24

OpenStudy (faiqraees):

@jhonyy9 2x-12=36 is an assumption

OpenStudy (legomyego180):

they are not equal

OpenStudy (faiqraees):

And a false assumption looking at the fact that both triangles have different values of length for perpendiculars

OpenStudy (llamasaremine):

i haven't done trigonometric ratios or whatever it is.

OpenStudy (llamasaremine):

that just confused me even more...

OpenStudy (anonymous):

not this side with one line y and write sin36 = 6/y so in this way you can calculi y and bc. you can use this in the other part too with side 5 and here on the left part this y will be y too so you can using sin(2x -12) = 5/y in this way you can getting the value of 2x-12 hope you understand me

OpenStudy (faiqraees):

@jhonyy9 I just pondered on the fact that you cant apply trigonometric ratios since it isn't a right angled triangle

OpenStudy (anonymous):

so than i think here is something mistaque in the text of this exercise - or not ?

OpenStudy (faiqraees):

Yeah there's lack of information

OpenStudy (llamasaremine):

thats all that is in the question (Sorry my computer was lagging)

OpenStudy (faiqraees):

Which chapter is this question related to?

OpenStudy (anonymous):

what is your opinion about these ? 36° ---- 6 2x-12 --- 5 --------------- (2x-12)*6 = 5*36 12x -72 = 180 12x = 252 x=252/12 x = 21

OpenStudy (legomyego180):

there isnt a lack of information

OpenStudy (llamasaremine):

lesson 6 unit 7 Triangle review (Message bored review not in lesson review.)

OpenStudy (faiqraees):

what is your opinion about these ? 36° ---- 6 2x-12 --- 5 --------------- (2x-12)*6 = 5*36 (The angles are not in ratio, this is a wrong equality) 12x -72 = 180 12x = 252 x=252/12 x = 21

OpenStudy (anonymous):

bc. these other two sides are congruent so in this way i think you can using this

OpenStudy (faiqraees):

That's only applicable when both triangles are congruent. One side is not same so isn't applicable

OpenStudy (anonymous):

so than how you wan solve it , in what way ?

OpenStudy (faiqraees):

I think there is lack of information.

OpenStudy (anonymous):

@ganeshie8 please

OpenStudy (anonymous):

@mathstudent55 @welshfella @zepdrix your opinion in this case ? ty

OpenStudy (welshfella):

i think all we can say is 2x - 12 < 36

OpenStudy (welshfella):

x < 24

OpenStudy (welshfella):

the 2 triangles are not congruent.